OpenStudy (anastasiagia):

Evaluate 32 + (6 − 2) ⋅ 4 − 6 over 3.

OpenStudy (legomyego180):

The answer is 14. I will post how I got this in just a sec.

OpenStudy (anastasiagia):

14 isnt one of the answers

OpenStudy (anastasiagia):

6 over three is a fraction

OpenStudy (legomyego180):

32+(6-2)*4-6\[\frac{ 32+(6-2)*4 }{ 3 }\] So remembering order of operations is key to solving this problem. Remember: Parentheeses Exponets Multiplication Division Addition Subtraction (Please Excuse My Dear Aunt Sally) So lets start with the parentheeses first. 6-2=4 So that gives us: 32+(6-2)*4-6\[\frac{ 32+4*4 }{ 3 }\] Now we can do our multiplication 4*4 is 16 so we have 32+(6-2)*4-6\[\frac{ 32+16 }{ 3 }\] Now we add the top together to get \[\frac{ 48}{ 3 }\] The answer is 16

OpenStudy (legomyego180):

Sorry my first answer was incorrect, I rushed through it

OpenStudy (anastasiagia):

16 isnt one of the options either

OpenStudy (anastasiagia):

20 23 50 26

OpenStudy (anastasiagia):

either one of those^

OpenStudy (legomyego180):

Im sorry your right, you need to factor

OpenStudy (anastasiagia):

?

OpenStudy (legomyego180):

Have you learned factoring?

OpenStudy (anastasiagia):

is it 20

OpenStudy (anastasiagia):

nope

OpenStudy (anastasiagia):

?

OpenStudy (legomyego180):

Could you post a pic of the problem?

OpenStudy (anastasiagia):

http://openstudy.com/study#

OpenStudy (anastasiagia):

helpppppppppppp

OpenStudy (legomyego180):

Sorry, I was picturing the problem wrong. \[3^2+(6-2)*4-\frac{ 6 }{ 3 }\] Now remember order of operations, parentheses and exponets first... \[9+4*4-\frac{ 6 }{ 3 }\] Now we move on to multiplication and division: six divided by three is two so we can rewrite this as: \[9+4*4-2\] You may be tempted to subtract 4-2 at this point but remember your order of operations, we have to multiply 4 times 4 first. \[9+16-2\] Now we just add and subtract and we have our answer, 9+16-2= 23 23 is the answer Sorry I took so long I misread the problem like 4 times.

OpenStudy (anastasiagia):

thank you so much